# Robust Clustering on High-Dimensional Data with Stochastic Quantization

### v1

@@ 1 Introduction
This paper explores advanced modifications of the Stochastic Quantization algorithm, incorporating accelerated variants of SGD [36, 44, 55] and adaptive learning rate techniques [12, 25, 53] to enhance convergence speed. Norkin et al. [41] provide a comprehensive comparison of various SGD variants, highlighting their respective advantages and limitations, while also offering convergence speed estimations.
@@ 2 Stochastic Quantization
Unlike traditional clustering methods that minimize the distance between each element of $\Xi=\{\xi_{i},i=1,\ldots,I\}$ and the nearest center $Y=\{y_{k},k=1,\ldots,K\}$, Stochastic Quantization conceptualizes the feature set $\Xi$ and cluster centers $Y$ as discrete probability distributions. The Wasserstein (or Kantorovich–Rubinstein) distance is employed to minimize distortion between these distributions when representing a compact distribution by a discrete one [29, 30]. Subsequent research [29, 40] has explored the application of quantization algorithms to solve optimal allocation problems for service centers, where each atom of the discrete distribution represents the location of facilities and customers, respectively.
@@
For a continuous probability distribution $\Xi$, we can interpret the objective function (3) as an unconstrained smooth stochastic optimization problem [14, 37, 41]:
$\min_{y=\{y_{1},\ldots,y_{K}\}\in Y^{K}\subset\mathbb{R}^{nK}}F(y_{1},\ldots,y_{k})=\mathbb{E}\tilde{F}(y,\xi)=\int_{\xi\in\Xi}\tilde{F}(y,\xi)P(d\xi)$ (6)
Assuming that the objective function $F(y)$ in (6) is differentiable and bounded from below, i.e., $\inf_{y\in Y}F(y)>-\infty$, we obtain the continuous representation of (4):
$F(y)=F(y_{1},\ldots,y_{k})=\mathbb{E}_{\xi}\min_{1\leq k\leq K}d(\xi_{i},y_{k})^{r}=\int\min_{1\leq k\leq K}d(\xi_{i},y_{k})^{r}P(d\xi)$ (7)
where the random variable $\xi$ may have a multimodal continuous distribution. The empirical approximation is:
@@
where $\{\xi_{i},i=1,\ldots,N\}$ are independent, identically distributed initial samples of the random variable $\xi$. If $K=1$ and $Y$ is convex, then problem (3) is unimodal and reduces to a convex stochastic optimization problem:
$\min_{y\in Y}[F(y)=\mathbb{E}_{\tilde{i}\sim p}d(\xi_{\tilde{i}},y)]$ (9)
However, for $K\geq 2$, the function $f(\xi_{i},y)=\min_{1\leq k\leq K}d(\xi_{i},y_{k}),y=(y_{1},\ldots,y_{K})$ is non-smooth and non-convex. In terms of [34, 39], if $f(\xi,y)$ is a random generalized differentiable function, its generalized gradient set can be calculated by the chain rule:
@@
The expected value function (4) is also generalized differentiable, and the set $\mathbb{E}_{\xi}\partial f(\xi,y)$ is a generalized gradient set of the function [34, 39]. Vectors $g(\xi)=(0,\ldots,0,g_{k},0,\ldots,0),k\in S(\xi,y),0\in\mathbb{R}^{n}$ are stochastic generalized gradients of the function $F(y_{1},\ldots,y_{K})$.
@@
4:     $\tilde{\xi}\in\{\xi_{i},\quad i=1,\ldots,I\}$triangleright Sample element from feature set
@@
6:     $g_{k}^{(t)}=r\|\tilde{\xi}-y_{k}^{(t)}\|^{r-2}(y_{k}^{(t)}-\tilde{\xi})$triangleright Calculate gradient
7:     $y_{k}^{(t)}:=\pi_{Y}(y_{k}^{(t)}-\rho g_{k}^{(t)})$triangleright Update nearest center
@@
$y_{k+1}=\pi_{Y}(y_{k}-\rho_{k}g_{k^{*}}),\quad\pi_{Y}(x)=\operatorname*{arg\,min}_{y\in Y}\|x-y\|,\quad y^{0}\in Y,\quad k\in\mathbb{N}$ (11)
@@
$\sum_{i=1}^{I}p_{i}\min_{1\leq k\leq K}d(\xi_{i},\pi_{Y_{k}}(\xi_{i}))^{r}$$\leq$ $\sum_{i=1}^{I}p_{i}\min_{y\in Y^{K}}\min_{1\leq k\leq K}d(\xi_{i},y_{k})^{r}$ (12)
$\leq$ $\min_{\{y_{k}\in Y_{k}\}}F(y_{1},\ldots,y_{K})$
The local convergence conditions of the stochastic generalized gradient method for solving problem (3) are determined in Theorem 2.1, with the proof provided in the works of [15] and [16].
@@
$g_{k}^{(t)}$ $=r\|\tilde{\xi}-y_{k}^{(t)}\|^{r-2}(y_{k}^{(t)}-\tilde{\xi})$ $k$$\in\{1,\ldots,K\}$ (13)
@@ 2.1 Adaptive Stochastic Quantization
Nesterov [36, 55] further improved the modified recurrent sequence (16) by introducing an extrapolation step for parameter estimation:
@@
Although modifications (16) and (17) can improve convergence speed, they often encounter the vanishing gradient problem on sparse data [8]. The root cause is the fixed learning rate value, which performs equal updates for both significant and insignificant model parameters. Duchi et al. [13] address this issue by introducing an adaptive learning rate $\tilde{\rho}_{k}=\rho_{k}/\sqrt{G_{k}+\varepsilon}$, where the hyperparameter value is normalized over the accumulated gradient value to increase the update for more significant parameters:
@@
where $G_{k}=G_{k-1}+g_{k^{*}}^{2}$ is a linear combination of accumulated gradients from previous iterations, and $\varepsilon\ll 10^{-8}$ is a denominator smoothing term. While approach (18) solves the convergence issue on sparse data, it introduces the problem of uncontrollable vanishing of the learning rate with each iteration, i.e., $\lim_{k\to\infty}\|\tilde{\rho}_{k}\|=0$. Tieleman et al. [53] proposed another approach for accumulated gradient normalization using a moving average $G_{k}=\beta G_{k-1}+(1-\beta)g_{k^{*}}^{2}$, which substitutes the denominator $G_{k}$ with a stochastic approximation of the expected value $\mathbb{E}G_{k}$ to control learning rate vanishing with an averaging multiplier $10<\beta<$ 1.
Kingma et al. [25] introduced a further modification to (18) by adding adaptive estimation of the gradient value $g_{k^{*}}$:
@@
Norkin et al. [41] provide a comprehensive overview of these adaptive parameter update strategies and present a detailed comparison of their convergence speed in various problem settings.
@@ 3 Traditional Clustering Model
The optimal clustering problem seeks to find K cluster centers $\{y_{1},...,y_{K}\}$ that minimize the sum of distances from each point $\xi$ to the nearest center. Lloyd’s KMeans algorithm [33] is a prominent method for solving quantization and clustering problems, with numerous extensions [22]. Bottou and Bengio [6] interpreted the KMeans algorithm as an analogue of Newton’s method and proposed several stochastic gradient descent algorithms for optimal clustering. These stochastic KMeans algorithms use only a subset of possible $\xi$ values or even a single element at each iteration.
@@
[33]. KMean iterative algorithm starts with the set of current cluster centers $\{y_{k}^{t},\>k=1,...,K\}$, the feature set $\{\xi_{i},\>i=1,...,I\}$ is subdivided into $K$ non-intersecting groups $\{I_{1}^{t},...,I_{K}^{t}\}$: point $\xi_{i}$ belongs to group $I_{s}^{t}$ if
$||\xi_{i}-y_{s}^{t}||=\min_{1\leq k\leq K}\{||\xi_{i}-y_{k}^{t}||\}$ (21)
Denote $N_{k}^{t}$ the number of points in group $I_{k}^{t}$, $N_{k}^{t}=0$ if $I_{k}^{t}=\emptyset$, $\sum_{k=1}^{K}N_{k}^{t}=1$. Remark that $I_{k}^{t}$ and $N_{k}^{t}$ depend on $y^{t}$. KMeans iteratively evaluates next cluster centers $y^{t+1}$ with the estimation:
@@
In [6] the form of KMeans algorithm (23) was connected to Newton’s step for solving at each iteration the smooth quadratic problem
$F^{t}(y)=\frac{1}{2}\sum_{k=1}^{K}\sum_{i\in I_{k}^{t}}||\xi_{i}-y_{k}||^{2}$ (24)
with block diagonal Hessian and with diagonal number $\frac{1}{N_{k}^{t}}$ in block $k$. In that paper it was also considered stochastic batch and online version of the stochastic gradient methods with learning rate $\frac{1}{t+1}$ for solving a sequence of problems (24).
The initial positions of cluster centers $\{y_{k}^{0},\>k=1,...,K\}$ are set either at random among $\{\xi_{i},\>i=1,...,I\}$ or according the algorithm KMeans++ [3, 38]. With KMeans++ initialization strategy, the rate of convergence to local optimum is estimated to be $\mathbb{E}[F]\leq 8(\ln k+2)F^{*}$, where $F^{*}$ an optimal solution [3]. Assume $\{y_{1}^{0},...,y_{k}^{0}\}\>(k<K)$ initial cluster centers have already been chosen. The next center $y_{k+1}^{0}\>(k+1<K)$ is sampled from the set $\{\xi_{i},\>i=1,...,I\}$ with probabilities:
@@
The next positions of the cluster centers $\{y_{1}^{t},...,y_{K}^{t}\}$ for $t>0$ are calculated as in the original Lloyd algorithm [33], the expectation-maximization (EM) approach to KMeans algorithm. Consider problem for $r=2$:
@@
Sculley [50] addresses the limitation of Lloyd algorithm with update rule (26), highlighting that objective function $F(y)$ calculation is expensive for large datasets, due to $O(K\cdot I\cdot d)$ time complexity for a given feature set $\{\xi_{i}\}$, where $d$ dimensionality of each sample $\xi_{i}$. Author proposed a solution by introducing a Mini-Batch KMeans modification to the Lloyd algorithm, where the set of points $\{\xi_{i},i=1,...,I\}$ is subdivided into non-intersecting subsets $I_{k}$, $k=1,...,K$, such that $i\in I_{k}$ if $||y-\xi_{i}||=\min_{k\in\{1,...,K\}}||y_{k}-\xi_{i}||$. Some $I_{k}$ may occur empty, for example, if $\max_{1\leq i\leq I}\min_{1\leq k\leq K}\text{dist}(y_{k},\xi_{i})<\min_{1\leq i\leq I}\text{dist}(y_{k},\xi_{i})$. In this modifications the generalized gradients of function is $g(y)=(g_{1}(y),...,g_{K}(y))$, where:
@@
Recent studies [52, 60] have examined stochastic KMeans algorithms as methods for solving corresponding non-convex, non-smooth stochastic optimization problems. However, the convergence properties of these algorithms lack rigorous validation. Let $N_{k}^{t}$ is the number of elements in $I_{k}\neq\emptyset$ at iteration $t$. If we chose $\rho_{t}$ dependent on $k$, namely, $\rho_{tk}=0.5\frac{I}{N_{k}^{t}}$, then process (30) becomes identical to KMeans one (22). Here $\rho_{tk}\leq 0.5I$ and can be rather large that does not guarantee convergence of (22). A more general choice can be $\rho_{tk}=\rho_{t}\frac{I}{N_{k}^{t}}$ with $\rho_{t}\geq 0$ satisfying conditions and thus $\lim_{t\to\infty}\max_{k}\rho_{tk}=0$.
3.1 Modifications of KMeans algorithm
One can consider a complement to the sequence (30) by the Cesaro trajectory averaging:
@@ 3.1 Modifications of KMeans algorithm
Conditions of convergence for this averaged sequence were studied in [34], in particular, they admit learning rate $\rho_{t}$ proportional to $1/\sqrt{t+1}$. A similar approach for KMeans generated sequences (22) aims to average sequence by the feature set size $N_{k}$:
$\tilde{y}_{k}^{t+1}=(1-\tilde{\sigma}_{t+1})\tilde{y}_{k}^{t}+\tilde{\sigma}_{t+1}y_{k}^{t+1},\>\tilde{\sigma}_{k,t+1}=\frac{1}{N_{k}^{t}}/\sum_{s=1}^{t+1}\frac{1}{N_{k}^{s}},\>k=1,...,K$ (32)
The standard KMeans algorithm requires finding the nearest cluster center $\operatorname*{arg\,min}_{1\leq k\leq K}||y_{k}-\xi_{i}||$ to each point $\{\xi_{i},\>i=1,...,I\}$. This can be a time consuming operation in case of very large number $I$. Moreover, points $\xi_{i}$ may be sampled sequentially from some continuous distribution and thus the sample $\{\xi_{i},\>i=1,2,...\}$ can be potentially arbitrary large. The Stochastic Quantization algorithm (31) uses only one sampled point $\tilde{\xi}^{t}$ at each iteration and thus only one problem $\operatorname*{arg\,min}_{1\leq k\leq K}||y_{k}-\tilde{\xi}^{t}||$ is solved at iteration $t$. But one can use a batch of $m$ such points $\{\tilde{\xi}^{t}_{i_{1}^{t}},...,\tilde{\xi}^{t}_{i_{m}^{t}}\}$ instead of the whole sample $\{\xi_{i}\}$, $m<I$.

### v2

@@ 1 Introduction
This paper explores advanced modifications of the Stochastic Quantization algorithm, incorporating accelerated variants of SGD [35, 44, 55] and adaptive learning rate techniques [24, 53] to enhance convergence speed. Norkin et al. [41] provide a comprehensive comparison of various SGD variants, highlighting their respective advantages and limitations, while also offering convergence speed estimations.
@@ 2 Stochastic Quantization
Unlike traditional clustering methods that minimize the distance between each element of $\Xi=\{\xi_{i},i=1,\ldots,I\}$ and the nearest center $Y=\{y_{k},k=1,\ldots,K\}$, Stochastic Quantization conceptualizes the feature set $\Xi$ and cluster centers $Y$ as discrete probability distributions. The Wasserstein (or Kantorovich–Rubinstein) distance is employed to minimize distortion between these distributions when representing a continuous distribution by a discrete one [28, 29]. Subsequent research [28, 40] has explored the application of quantization algorithms to solve optimal allocation problems for service centers, where each atom of the discrete distribution represents the location of facilities and customers, respectively.
@@
For a continuous probability distribution $P(d\xi)$, we can interpret the objective function (3) as a mathematical expectation in a stochastic optimization problem [13, 36, 41]:
$\min_{y=\{y_{1},\ldots,y_{K}\}\in Y^{K}\subset\mathbb{R}^{nK}}\left[F(y_{1},\ldots,y_{k})=\mathbb{E}f(y,\xi)=\int_{\xi\in\Xi}f(y,\xi)P(d\xi)\right]$ (6)
with
$f(y,\xi)=\min_{1\leq k\leq K}d(\xi,y_{k})^{r},$ (7)
where the random variable $\xi$ may have a multimodal continuous distribution. The empirical approximation of $F(y)$ in (6) is:
@@
where $\{\xi_{i},i=1,\ldots,N\}$ are independent, identically distributed initial samples of the random variable $\xi$. If $K=1$, $Y$ is convex, and $r\geq 1$, then problem (3) is unimodal and reduces to a convex stochastic optimization problem:
$\min_{y\in Y}[F(y)=\mathbb{E}_{\tilde{i}\sim p}d(\xi_{\tilde{i}},y)^{r}]$ (9)
However, for $K\geq 2$, the function $f(\xi,y)=\min_{1\leq k\leq K}d(\xi,y_{k})^{r},y=(y_{1},\ldots,y_{K})$ is non-smooth and non-convex. In terms of [33, 38], $f(\xi,y)$ is a random generalized differentiable function, its generalized gradient set can be calculated by the chain rule:
@@
The expected value function (4) is also generalized differentiable, and the set $\mathbb{E}_{\xi}\partial f(\xi,y)$ is a generalized gradient set of the function $F$ [33, 38]. Vectors $g(\xi)=(0,\ldots,0,g_{k},0,\ldots,0),k\in S(\xi,y),0\in\mathbb{R}^{n}$, are stochastic generalized gradients of the function $F(y_{1},\ldots,y_{K})$.
@@
4:     $\tilde{\xi}\in\{\xi_{i},\quad i=1,\ldots,I\}$triangleright Sample an element from the feature set
@@
6:     $g_{k}^{(t)}=r\|\tilde{\xi}-y_{k}^{(t)}\|^{r-2}(y_{k}^{(t)}-\tilde{\xi})$triangleright Calculate a gradient
7:     $y_{k}^{(t)}:=\pi_{Y}(y_{k}^{(t)}-\rho g_{k}^{(t)})$triangleright Update the nearest center
@@
$y_{k+1}=\pi_{Y}(y_{k}-\rho_{k}g_{k^{*}}),\quad\pi_{Y}(x)=\operatorname*{arg\,min}_{y\in Y}\|x-y\|,\quad y^{0}\in Y,\quad k\in\mathbb{N},$ (11)
@@
$\min_{\{y_{k}\in Y_{k}\}}F(y_{1},\ldots,y_{K})$$\geq$ $\sum_{i=1}^{I}p_{i}\min_{y\in Y_{1}\times\ldots\times Y_{K}}\min_{1\leq k\leq K}d(\xi_{i},y_{k})^{r}$ (12)
$\geq$ $\sum_{i=1}^{I}p_{i}\min_{1\leq k\leq K}d(\xi_{i},\pi_{k}(\xi_{i})^{r}.$
The local convergence conditions of the stochastic generalized gradient method for solving problem (3) are determined in Theorem 2.1, with the proof provided in [14, 15].
@@
$g_{k}^{(t)}$ $=r\|\tilde{\xi}^{(t)}-y_{k}^{(t)}\|^{r-2}(y_{k}^{(t)}-\tilde{\xi}^{(t)})$ $k$$\in\{1,\ldots,K\}$ (13)
@@ 2.1 Adaptive Stochastic Quantization
Nesterov [35, 55] further improved the modified recurrent sequence (16) by introducing an extrapolation step for parameter estimation (Nesterov Accelerated Gradient or NAG):
@@
Although modifications (16) and (17) can improve convergence speed, they often encounter the vanishing gradient problem on sparse data [8]. The root cause is the fixed learning rate value, which performs equal updates for both significant and insignificant model parameters. Duchi et al. [12] address this issue by introducing an adaptive learning rate $\tilde{\rho}_{k}=\rho_{k}/\sqrt{G_{k}+\varepsilon}$, where the hyperparameter value is normalized over the accumulated gradient value to increase the update for more significant parameters (AdaGrad):
@@
where $G_{k}=G_{k-1}+g_{k^{*}}^{2}$ is a linear combination of accumulated gradients from previous iterations, and $\varepsilon\ll 10^{-8}$ is a denominator smoothing term. While approach (18) solves the convergence issue on sparse data, it introduces the problem of uncontrollable vanishing of the learning rate with each iteration, i.e., $\lim_{k\to\infty}|\tilde{\rho}_{k}|=0$. Tieleman et al. [53] proposed another approach (RMSProp) for accumulated gradient normalization using a moving average $G_{k}=\beta G_{k-1}+(1-\beta)g_{k^{*}}^{2}$, which substitutes the denominator $G_{k}$ with a stochastic approximation of the expected value $\mathbb{E}G_{k}$ to control learning rate vanishing with an averaging multiplier $10<\beta<$ 1.
Kingma et al. [24] introduced a further modification to (18) by adding adaptive estimation of the gradient value $g_{k^{*}}$ (ADAM):
@@
Norkin et al. [41] provide an overview of these adaptive parameter update strategies and present a detailed comparison of their convergence speed in various problem settings.
@@ 3 Traditional Clustering Model
The optimal clustering problem seeks to find $K$ cluster centers $\{y_{1},...,y_{K}\}$ that minimize the sum of distances from each point $\xi$ to the nearest center. Lloyd’s K-Means algorithm [32] is a prominent method for solving quantization and clustering problems, with numerous extensions [21]. Bottou and Bengio [6] interpreted the K-Means algorithm as an analogue of Newton’s method and proposed several stochastic gradient descent algorithms for optimal clustering. These stochastic K-Means algorithms use only a subset of possible $\xi$ values or even a single element at each iteration.
@@
[32]. K-Means iterative algorithm starts with the set of current cluster centers $\{y_{k}^{t},\>k=1,...,K\}$, the feature set $\{\xi_{i},\>i=1,...,I\}$ is subdivided into $K$ non-intersecting groups $\{I_{1}^{t},...,I_{K}^{t}\}$: point $\xi_{i}$ belongs to group $I_{s}^{t}$ if
$\|\xi_{i}-y_{s}^{t}\|^{2}=\min_{1\leq k\leq K}\{\|\xi_{i}-y_{k}^{t}\|^{2}\},\;\;\;r>0.$ (21)
Denote $N_{k}^{t}$ the number of points in group $I_{k}^{t}$, $N_{k}^{t}=0$ if $I_{k}^{t}=\emptyset$, $\sum_{k=1}^{K}N_{k}^{t}=I$. Remark that $I_{k}^{t}$ and $N_{k}^{t}$ depend on $y^{t}$. K-Means iteratively evaluates next cluster centers $y^{t+1}$ with the estimation:
@@
In [6] the form of K-Means algorithm (23) was connected to Newton’s step for solving at each iteration the smooth quadratic problem
$\min_{y_{1},\ldots,y_{K}}\left[F^{t}(y)=\frac{1}{2}\sum_{k=1}^{K}\sum_{i\in I_{k}^{t}}||\xi_{i}-y_{k}||^{2}\right].$ (24)
with block diagonal Hessian and with diagonal number $\frac{1}{N_{k}^{t}}$ in block $k$. Moreover, it is easy to see that (23) is the exact analytical solution of the unconstrained quadratic minimization problem (24) under fixed partition $\{I_{1}^{t},\ldots,I_{K}^{t}\}$ of the index set $\{1,\ldots,I\}$. In that paper it was also considered stochastic batch and online version of the stochastic gradient methods with learning rate $\frac{1}{t+1}$ for solving a sequence of problems (24) but without rigorous convergence analysis.
The initial positions of cluster centers $\{y_{k}^{0},\>k=1,...,K\}$ are set either at random among $\{\xi_{i},\>i=1,...,I\}$ or according the algorithm K-Means++ [3, 37]. With K-Means++ initialization strategy, the rate of convergence to local optimum is estimated to be $\mathbb{E}[F]\leq 8(\ln k+2)F^{*}$, where $F^{*}$ an optimal solution [3]. Assume $\{y_{1}^{0},...,y_{k}^{0}\}\>(k<K)$ initial cluster centers have already been chosen. The next center $y_{k+1}^{0}\>(k+1<K)$ is sampled from the set $\{\xi_{i},\>i=1,...,I\}$ with probabilities:
@@
The next positions of the cluster centers $\{y_{1}^{t},...,y_{K}^{t}\}$ for $t>0$ are calculated as in the original Lloyd algorithm [32], the expectation-maximization (EM) approach to K-Means algorithm. Consider problem for $r=2$:
@@
Sculley [50] addresses the limitation of Lloyd algorithm with update rule (26), highlighting that objective function $F(y)$ calculation is expensive for large datasets, due to $O(K\cdot I\cdot d)$ time complexity for a given feature set $\{\xi_{i}\}$, where $d$ dimensionality of each sample $\xi_{i}$. The author proposed a solution by introducing a Mini-Batch K-Means modification to the Lloyd algorithm, where the set of points $\{\xi_{i},i=1,...,I\}$ is subdivided into non-intersecting subsets $I_{k}$, $k=1,...,K$, such that $i\in I_{k}$ if $\|y_{k}-\xi_{i}\|=\min_{k^{\prime}\in\{1,...,K\}}\|y_{k^{\prime}}-\xi_{i}\|$. Some $I_{k}$ may occur empty, for example, if
$\max_{1\leq i\leq I}\min_{1\leq k\leq K}\|y_{k}-\xi_{i}\|<\min_{1\leq i\leq I}\|y_{k}-\xi_{i}\|.$
In this modifications the generalized gradients of function $F(y)$ is $g(y)=(g_{1}(y),...,g_{K}(y))$, where:
@@
Recent studies [52, 60] have examined stochastic K-Means algorithms as methods for solving corresponding non-convex, non-smooth stochastic optimization problems. However, the convergence properties of these algorithms lack rigorous validation. Let $N_{k}^{t}$ is the number of elements in $I_{k}\neq\emptyset$ at iteration $t$. If we chose $\rho_{t}$ dependent on $k$, namely, $\rho_{tk}=0.5\frac{I}{N_{k}^{t}}$, then process (30) becomes identical to K-Means one (22). Here $\rho_{tk}\leq 0.5I$ and can be rather large that does not guarantee convergence of (22). A more general choice can be $\rho_{tk}=\rho_{t}\frac{I}{N_{k}^{t}}$ with $\rho_{t}\geq 0$ satisfying conditions (14) and thus $\lim_{t\to\infty}\max_{k}\rho_{tk}=0$.
3.1 Modifications of K-Means algorithm
Robust clustering model assumes solving problem (3), (4) with parameter $r<2$. Such choice of parameter $r$ makes the quantization and clustering model more robust to outliers. However, then stochastic generalized gradients of the objective function should be calculated by formula (2) and the stochastic clustering algorithm takes form (11).
One also can consider a complement to the sequence (30) by the Cesàro trajectory averaging:
@@ 3.1 Modifications of K-Means algorithm
Conditions of convergence for this averaged sequence were studied in [33], in particular, they admit learning rate $\rho_{t}$ proportional to $1/\sqrt{t+1}$. A similar approach for K-Means generated sequences (22) aims to average sequence by the feature set size $N_{k}$:
$\tilde{y}_{k}^{t+1}=(1-\tilde{\sigma}_{t+1})\tilde{y}_{k}^{t}+\tilde{\sigma}_{t+1}y_{k}^{t+1},\>\tilde{\sigma}_{k,t+1}=\frac{1}{N_{k}^{t+1}}/\sum_{s=1}^{t+1}\frac{1}{N_{k}^{s}},\>k=1,...,K$ (32)
The standard K-Means algorithm requires finding the nearest cluster center $\operatorname*{arg\,min}_{1\leq k\leq K}\|y_{k}-\xi_{i}\|$ to each point $\{\xi_{i},\>i=1,...,I\}$. This can be a time consuming operation in case of very large number $I$. Moreover, points $\xi_{i}$ may be sampled sequentially from some continuous distribution and thus the sample $\{\xi_{i},\>i=1,2,...\}$ can be potentially arbitrary large. The Stochastic Quantization algorithm (31) uses only one sampled point $\tilde{\xi}^{t}$ at each iteration and thus only one problem $\operatorname*{arg\,min}_{1\leq k\leq K}\|y_{k}-\tilde{\xi}^{t}\|$ is solved at iteration $t$. But one can use a batch of $m$ such points $\{\tilde{\xi}^{t}_{i_{1}^{t}},...,\tilde{\xi}^{t}_{i_{m}^{t}}\}$ instead of the whole sample $\{\xi_{i}\}$, $m<I$.